OpenStudy (anonymous):

Hey, can someone explain an equation to me step by step? Ill fan and give a Medal. I have to solve this using substitution, and I'm having a bit of difficulty. y = 9x + 14 29x +7y = 6 I know the answer is ( -1, 5) Im just having difficulty understanding how you get that answer

OpenStudy (anonymous):

since they gave you y= ... start by plugging that in for y in the equation below that

OpenStudy (anonymous):

\[29 x +7 (9x+14)=6\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Wouldn't is be 7*14?

OpenStudy (anonymous):

\[29x + 63x + 98= 6\]

OpenStudy (anonymous):

sorry, got mixed up for a sec

OpenStudy (anonymous):

Its fine :) so you add 29 and 63 equaling 92x

OpenStudy (anonymous):

\[92x= 6-98\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[x= -92/92 = -1\]

OpenStudy (anonymous):

Then plug in x and solve for y right? Thanks for the help!

OpenStudy (anonymous):

now that you have solved for "x". plug that same value in for x to get y

OpenStudy (anonymous):

Sorry, Im having technical difficulties so Im replying slowly.

OpenStudy (anonymous):

\[y= 9(-1) +14 = 14-9 = 5\]

OpenStudy (anonymous):

np